# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v4

@@ 2.2 Optimization
2.3 Forecast
The target variable for forecasting is the aggregate uncontrollable load of the energy community in Figure 2. Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model. The gradient boosting implementation is LightGBM [29]. Predictions are made directly for the next 24 hours with 24 independent models. The scenario sets are generated with added gaussian noise. The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al. [27] and the code repository of the winner solution is available on GitLab [30]. The evaluation conducted in this study is available in the code repository [31].
While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings. This setup follows the problem setting given in the Citylearn Challenge 2022 [27]. Although this train-test splitting is not typical in applied tasks, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
@@ 2.3.1 Rolling Origin Window
In practical applications, the forecast is issued periodically for a finite number of future time steps. This process is run iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps. A schematic of this process is shown in Figure 3. This approach involves continuously moving the time window forward by a certain step size after each forecasting iteration. For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
Figure 3: Updating Rolling Origin Window. The forecast
is updated every $v_{f}$ time steps and possible overlaps between prediction windows are shown.
The rolling horizon setup enables an update of the prediction using the most recent information. When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions. The resulting overlapping interval is highlighted in light grey in Figure 3. These overlaps are considered when evaluating stability of the forecast.
@@ 2.3.2 Stability
Stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed in [20], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 4.
@@
Figure 4: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions. The vertical stability is the difference between the predictions produced at different origins for the same time period. If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal. Horizontal stability characterizes the difference between predictions generated at the same origin. The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
Similarly to the evaluation in [32], a stability metric is measured for point forecasts using the mean absolute change (MAC). The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the change of predictions within the horizon window generated at the same origin — ${MAC}_{H}$
@@
$\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (4)
@@
$\operatorname{MAC}_{H}=\frac{\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|}{H-1}$ (5)
@@ 2.3.3 Stability of Probabilistic Forecasts
To assess stability across time, SDC evaluates both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates). For $N$ number of scenarios, these are defined as:
@@
$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (6)
Vertical Stability (SDCV):
$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (7)
2.3.4 Accuracy of Determinstic Forecasts

### v5

@@ 2.2.1 Rolling Origin Window
In practical applications, forecasts are issued periodically for a finite number of future time steps. This process runs iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps. Figure 2 illustrates this approach, which involves continuously moving the time window forward by a certain step size after each forecasting iteration.
  Input Data
 $\rightarrow$ Forecast Horizon
$t_{0}$$t_{1}$$t_{2}$$t_{3}$$t_{4}$$t_{5}$$t_{6}$$t_{7}$$t_{8}$orecast1Forecast2Forecast3Forecast4Forecast Origins
Figure 2: Rolling Origin Forecasting with Fixed Input Window and Forecast Horizon. The diagram illustrates how forecasts with horizon $H=2$ are updated every $v_{f}=1$ time steps, creating overlaps between prediction windows.
The rolling horizon setup enables prediction updates using the most recent information. When the commitment period is shorter than the forecast horizon, regular updates create overlaps between prediction windows, as forecasts for a particular future time point are released in multiple revisions. These overlaps are considered when evaluating forecast stability.
@@ 2.2.2 Stability
Forecast stability is defined as the variability of predictions over time. Following the categorization proposed by Godahewa et al. [2023], we distinguish between vertical and horizontal stability, as illustrated in Figure 3.
@@
Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions.
Vertical stability refers to the consistency between predictions made at different origins for the same target time period. Horizontal stability, on the other hand, characterizes the variance of predictions within the same forecast window.
3 Theoretical analysis of performance bounds
This section presents an analysis of the performance bounds for Fixed Horizon Control (FHC) algorithms with commitment level $v$. We derive upper bounds on the competitive difference, the gap between FHC and the optimal clairvoyant policy. Our analysis reveals a fundamental trade-off between forecast accuracy and switching costs that depends critically on how frequently policies are updated. We first develop these bounds in a deterministic setting, then extend them to stochastic optimization where forecasts are represented as scenario ensembles. Finally, we establish the theoretical connection between forecast stability and switching costs, demonstrating why stable forecasts can significantly improve performance in environments with high switching costs.
We adopt the framework introduced by Chen et al. [2016] for analyzing online convex optimization (OCO) problems with switching costs and noisy predictions. In this setting, a decision-maker sequentially selects actions $x_{t}\in\mathcal{F}\subset\mathbb{R}^{n}$ over a time horizon $T$, facing unknown convex cost functions $h_{t}(x_{t})$ and incurring a switching cost $\beta\|S_{t}-S_{t-1}\|$.
$\operatorname{cost}(ALG)=\sum_{t=1}^{T}h_{t}(x_{t})+\sum_{t=1}^{T}\beta\|S_{t}-S_{t-1}\|$ (2)
The true state $y_{t}$ at each time $t$ is unknown at the time of decision-making and must be predicted. Forecasts $\hat{y}_{t|\tau}$ are generated at time $\tau<t$ and are corrupted by noise, modeled as:
$y_{t}-\hat{y}_{t|\tau}=\sum_{s=\tau+1}^{t}f(t-s)e(s),$ (3)
where $e(s)$ are i.i.d. zero-mean noise terms with covariance matrix $R_{e}$, and $f(\cdot)$ is a deterministic impulse response function capturing the correlation structure of the forecast errors.
The performance of an online algorithm is assessed using the competitive difference, defined as the expected cost gap between the online algorithm and the offline optimal (oracle) policy. Formally, an online algorithm ALG, is said to have a competitive difference of $\rho(T)$ if:
$\sup_{\hat{y}}\mathbb{E}_{e}\left[\operatorname{cost}(\text{ALG})-\operatorname{cost}(\text{OPT})\right]\leq\rho(T)$ (4)
where the expectation is taken with respect to the prediction noise sequence $\{e(t)\}_{t=1}^{T}$. The offline optimal policy OPT has full access to future outcomes and is thus not affected by forecast errors. In contrast, the performance of online algorithms like Fixed Horizon Control (FHC) or Receding Horizon Control (RHC) is directly impacted by the structure and magnitude of the forecast noise.
Following this framework, we derive bounds on the expected performance degradation of FHC-type algorithms as a function of the commitment level $v$, the forecast error correlation $\|f_{v}\|$, and the switching cost penalty $\beta$. As a starting point, we refer to Theorem 1 in Chen et al. [2016], and adapt it to our context.
$\operatorname{cost}(FHC(v))\leq\operatorname{cost}(OPT)+2M_{k}\beta D+2G\sum_{t=1}^{T}\|S_{t}-S_{t\mid\phi_{k}(t)}\|$ (5)
where $M$ is the number of updates, $k$ is the commitment level, $D$ is the maximum switching cost, and $G$ is the Lipschitz constant of the cost function
Here, in contrast to the original paper, we derive the expectation over the forecast horizon for the FHC algorithm with commitment level $v$:
$\mathbb{E}[\operatorname{cost}(FHC(v))]$ $\leq\operatorname{cost}(OPT)+2M_{k}\beta D+2G\sum_{t=1}^{T}\mathbb{E}\left[\|S_{t}-S_{t\mid t-k}\|\right]$ (6)
$\leq\operatorname{cost}(OPT)+2(T/v)\beta D+2GT\cdot\|f_{v}\|^{\alpha}$ (7)
as $M_{k}$ is the number of optimization updates, which is equal to $T/v$ for the FHC(v) algorithm.
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (8)
@@ 3 Theoretical analysis of performance bounds
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|S_{t}-S_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition. We assume $\alpha\geq 1$ to ensure that the cost function is convex.
3.1 Deterministic Optimization
In practical applications, forecast errors often exhibit exponentially decaying temporal correlations. We model this behavior using a decay parameter $a\in(0,1)$ and define the Frobenius norm of the correlation impulse response as:
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (10)
where $c$ is a constant scaling factor and $s$ is the time lag between the forecast target time and forecast issuance. We assume zero-mean i.i.d. noise vectors $e(s)$ with covariance $\mathbb{E}[e(s)e(s)^{T}]=R_{e}$ and trace $\operatorname{trace}(R_{e})=\sigma^{2}$. The cumulative forecast error norm up to commitment level $v$ becomes:
$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\|R_{e}^{1/2}\|_{F}^{2}\cdot\|f(s)\|_{F}^{2}$
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\cdot\frac{1-a^{2(v+1)}}{1-a^{2}}$
Taking the square root and using the Lipschitz continuity of $h$ with constant $G$, we have:
$\|f_{v}\|=c\sigma\sqrt{\frac{1-a^{2(v+1)}}{1-a^{2}}}$ (11)
Substituting into the expected cost bound for the FHC algorithm, we obtain:
$\mathbb{E}[\operatorname{cost}(FHC(v))]\leq\mathbb{E}[\operatorname{cost}(OPT)]+\frac{2T\beta D}{v}+2GT\|f_{v}\|$ (12)
Define:
$A$ $=2T\beta D$
$B$ $=\frac{2GTc\sigma}{\sqrt{1-a^{2}}}$
Then the bound becomes:
$\mathbb{E}[\operatorname{cost}(FHC(v))]\leq\mathbb{E}[\operatorname{cost}(OPT)]+\frac{A}{v}+B\cdot\sqrt{1-a^{2(v+1)}}$ (13)
This formulation clearly reflects the trade-off. The term $\frac{A}{v}$ is the cost of policy switching and decreases with larger commitment $v$. The term $B\cdot\sqrt{1-a^{2(v+1)}}$ captures forecast error accumulation and increases with $v$ since $a^{2(v+1)}\to 0$ as $v$ increases. The role of each parameter is summarized in Table 2. Therefore, the expected competitive difference exhibits a U-shaped curve with respect to $v$, where a minimal cost is achieved by balancing the cost of frequent switching and the degradation due to forecast staleness. This quantifies the intuition that updating too frequently incurs high switching costs, while updating too infrequently leads to performance loss due to outdated forecasts. This function run with different sets of parameters is demonstrated in Figure 7. It is observed that the low switching costs lead to a linear increase in the expected competitive difference, while the high switching costs lead to a formation of a U-shaped curve.
Table 2: Key parameters in performance bound
Symbol Represents Interpretation
$A$ Switching Cost Weight Scales with horizon $T$, penalty $\beta$, and max switching distance $D$. Larger $A$ increases the switching cost.
$B$ Forecast Error Sensitivity Incorporates cost weight $G$, horizon $T$, and noise parameters. Amplifies penalty for forecast inaccuracy.
$a$ Correlation Decay Rate Controls how quickly forecast errors become uncorrelated over time. Lower values of $a$ mean errors at future steps are largely independent.
Figure 4: Illustration of the FHC expected competitive difference as a function of the commitment level for (a) Low and High switching costs with predictions of exponentially decaying correlations, and (b) Stochastic and Deterministic optimizations. Stochastic scenarios exhibit a lower level of forecast error sensitivity ($B_{s}<B$), shifting the optimal commitment level to higher values.
3.2 Stochastic Optimization over Scenarios
We now analyze the performance of Fixed Horizon Control (FHC) under a stochastic optimization setting where forecasts are provided as a set of $n$ equiprobable scenarios, denoted $\{\xi_{i}\}_{i=1}^{n}$, each representing a possible trajectory of future values over the horizon $v$. For each scenario $\xi_{i}$, the FHC algorithm yields a policy $x^{(i)}=(x^{(i)}_{t})_{t=1}^{T}$. The overall cost is computed by averaging over the scenario outcomes:
@@ 3.2 Stochastic Optimization over Scenarios
$\mathbb{E}\left[\operatorname{cost}(FHC_{\text{stochastic}}(v))\right]=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\operatorname{cost}(FHC(v,\xi_{i}))\right].$ (14)
Let $f_{v}^{(i)}\in\mathbb{R}^{d}$ denote the cumulative forecast error over horizon $v$ under scenario $\xi_{i}$. The expected cost includes a term of the form $\|f_{v}^{(i)}\|^{\alpha}$, which contributes to the second term in the performance bound.
Assuming $\alpha\geq 1$, the function $x\mapsto\|x\|^{\alpha}$ is convex, and Jensen’s inequality yields:
$\left\|\frac{1}{n}\sum_{i=1}^{n}f_{v}^{(i)}\right\|^{\alpha}\leq\frac{1}{n}\sum_{i=1}^{n}\left\|f_{v}^{(i)}\right\|^{\alpha}.$ (15)
This inequality implies that the effective contribution of forecast uncertainty to the cost is lower in the stochastic case than in the deterministic case, since:
$2GT\left\|\mathbb{E}[f_{v}^{(i)}]\right\|^{\alpha}\leq 2GT\cdot\frac{1}{n}\sum_{i=1}^{n}\left\|f_{v}^{(i)}\right\|^{\alpha}.$ (16)
Thus, scenario averaging introduces a variance reduction effect, which improves robustness to forecast noise. This insight leads to the following bound:
$\mathbb{E}\left[\operatorname{cost}(FHC_{\text{stochastic}}(v))\right]\leq\mathbb{E}\left[\operatorname{cost}(OPT)\right]+\frac{A}{v}+B_{s}(1-a^{2(v+1)})^{1/2},$ (17)
where $B_{s}<B$ captures the effect of scenario averaging. This result highlights a key theoretical benefit of stochastic FHC: the aggregation over multiple forecast paths leads to a lower effective forecast variance and therefore lower forecast error sensitivity. It is shown in Figure 4(b), that the vertex of the U-shaped curve shifts to higher values of $v$ in the stochastic case. The fundamental trade-off between switching costs (decreasing with $v$) and forecast accuracy (typically decreasing with more frequent updates) persists, but the stochastic case enables a longer commitment horizon before the forecast error term dominates.
3.3 Role of Forecast Stability in Optimization Performance
The theoretical performance bounds in Equation (9) consist of two main components: (i) a term proportional to the switching cost, $\frac{2T\beta D}{v}$, and (ii) a forecast error penalty, $2GT\|f_{v}\|^{\alpha}$. While the second term captures the accuracy of the forecast over the horizon $v$, the first term reflects the cost incurred by changing policies across planning updates. In this section, we formalize the connection between the stability of forecasts and the incurred switching costs. Let $x_{t}$ denote the decision policy applied at time $t$, and let this policy be a deterministic function of the forecast $\hat{y}_{t}$, such that:
$x_{t}=\mathcal{M}(\hat{y}_{t}),$ (18)
where $\mathcal{M}:\mathbb{R}^{H}\to\mathbb{R}^{d}$ is a policy generation function (e.g., the solution to an optimization problem over horizon $H$), and is assumed to be Lipschitz continuous with constant $L_{\mathcal{M}}$:
$\|\mathcal{M}(\hat{y}_{t})-\mathcal{M}(\hat{y}_{t-v})\|\leq L_{\mathcal{M}}\cdot\|\hat{y}_{t}-\hat{y}_{t-v}\|.$ (19)
Using this, the switching cost term can be bounded as:
$\sum_{t=1}^{T/v}\beta\|S_{t}-S_{t-v}\|\leq\beta L_{\mathcal{M}}\sum_{t=1}^{T/v}\|\hat{y}_{t}-\hat{y}_{t-v}\|.$ (20)
The term $\|\hat{y}_{t}-\hat{y}_{t-v}\|$ represents the change in forecast between updates, which corresponds to the concept of vertical forecast stability—the consistency of predictions made at different time origins for the same target time. A common empirical proxy for this is the Mean Absolute Change (MAC) metric:
$\operatorname{MAC}_{V}=\frac{1}{H-1}\sum_{i=1}^{H-1}|\hat{y}_{t+i|t}-\hat{y}_{t+i|t-v}|.$ (21)
Thus, improving the vertical stability of forecasts directly reduces the magnitude of policy changes, and therefore the incurred switching costs. This establishes a theoretical link between empirical measures of forecast stability and the first term in the performance bound (9). The MAC metric and appropriate stability metrics for the stochastic case are discussed in detail in the next section.
4 Experimentation and Results
In this section, we transition from theoretical analysis to practical implementation and evaluation. While the theoretical model provides important insights, real-world energy management systems face additional complexities including non-trivial forecast error correlations, potential non-convexity in optimization problems, and the challenges of coordinating multiple distributed assets. We present empirical results from applying the proposed framework to a battery scheduling problem, evaluating how different commitment periods affect system performance under varying forecast qualities. First, the current section outlines the experimental setup and case study, followed by a detailed analysis of the results. The analysis focuses on the relationship between forecast accuracy, stability, and the impact of switching costs on the optimal decisions made in the battery scheduling process.
4.1 Application to Battery Scheduling in Energy Management Systems
The decision-maker must balance the benefits of plan revisions against the costs of switching. This balance is particularly critical in energy management systems, where rapid changes in operation can strain infrastructure, reduce equipment lifespan, and lead to additional expenses or inefficiencies. To investigate these trade-offs in a realistic setting, we apply the concepts of online optimization with switching costs to battery storage management within a multi-agent energy management system. Our case study is based on the CityLearn Challenge 2022 Nweye et al. [2022], which provides a comprehensive dataset from an actual grid-interactive community.
The central challenge involves multi-agent scheduling of energy storage across a one-year period with hourly granularity. Figure 5 illustrates the problem setup. The data of 17 buildings used in this study is derived from a real-world zero net energy community in Fontana, California, USA, previously studied for grid integration of zero net energy developments as part of the California Solar Initiative program Narayanamurthy et al. [2016]. Each building has a battery with specific charging/discharging characteristics, and the system must coordinate these resources efficiently. The objective is to develop control policies that minimize grid electricity costs, carbon emissions and ramping while increasing the load factor. The associated ramping costs constitute the switching costs in this context. The combined objective score with the grid cost is defined as:
[…]
$\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}},D\right)$
where $C$ are electricity costs, $G$ are carbon emissions, and $D$ is the grid score. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. Full details of the objective score and relevant constraints are provided in A.1.
The proposed framework was implemented using the CityLearn simulation environment, an open-source platform designed for evaluating energy management systems in grid-interactive communities [^1] https://github.com/intelligent-environments-lab/CityLearn. Originally created for reinforcement learning (RL), the platform now also supports alternative methods like Model Predictive Control (MPC) and rule-based control. It simulates a range of energy assets, including batteries, solar panels, heat pumps, and electric vehicles, all of which play a role in optimizing energy consumption across a grid-connected community.
Figure 5: System setup (a) and data flow (b) for the battery scheduling problem. The data flow diagram illustrates how input features are processed through forecasting models with varying update frequencies to generate load predictions, which then inform the optimization stage to produce battery control actions.
4.2 Implementations of Forecasting and Optimization
The system architecture consists of two primary components: forecasting and optimization, as shown in Figure 5(b). The forecasting module generates load predictions, which serve as inputs to the optimization model. Both forecasting and optimization follow a rolling horizon approach. While the commitment period $v$ is commonly the same for both stages, updates are not inherently required to be synchronized. Therefore, we distinguish between $v_{F}$ (forecast update frequency) and $v_{O}$ (optimization update frequency) when necessary.
For the forecasting task, we predict the aggregate uncontrollable load of the energy community. The training set comprises data of 5 buildings, with the remaining 12 buildings used for validation and testing. The forecasts are generated using a combination of gradient boosting regression trees (LightGBM Ke et al. [2017]) and linear least squares regression. The method follows the procedure of the winning solution, outlined in Nweye et al. [2022]. We train the forecasting model across all buildings, then apply it to 7 out-of-sample buildings for testing, a practice commonly used in scenarios with limited historical data but similar building characteristics, as discussed in Grabner et al. [2023]. The forecasts are generated for a 24-hour horizon, with 24 separate models predicting each hour. To account for forecast uncertainty, we generate multiple forecast scenarios by adding Gaussian noise to the base predictions.
The optimization model, implemented with Pyomo Hart et al. [2017] and solved using Gurobi Gurobi Optimization, LLC [2024], aims to minimize a composite cost function that incorporates electricity costs, ramping penalties, switching costs, and carbon emissions. The optimization problem is solved for the same 24-hour horizon as the forecasting model, subject to revision every $v_{O}$ steps. At the optimization stage, both deterministic and stochastic forecast settings are considered. In the deterministic case, a single forecast trajectory is used to determine the optimal control policy. In the stochastic case, multiple forecast scenarios are considered, and the optimization is performed over 75 scenarios, minimizing the expected cost. This allows the system to better handle forecast uncertainty and improves robustness by optimizing expected performance across scenarios.
The optimization and forecasting steps are coordinated through the CityLearn platform, which integrates data from multiple sources, including weather forecasts, building characteristics, and historical consumption patterns. The resulting control policies, which dictate how the batteries should charge and discharge, are evaluated based on their ability to reduce grid electricity costs, minimize ramping costs, and lower carbon emissions. The implementation is available on GitHub [^2] https://github.com/ujohn33/Predict-Optimize-Revise and is fully reproducible, with optimization logs available upon request. The detailed mathematical formulation of the optimization problem and the constraints imposed on the system are provided in A.2.
4.3 Evaluation metrics
Similarly to the evaluation in [Van Belle et al., 2023], a stability metric is measured for point forecasts using the mean absolute change (MAC). The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the variation of predictions within the horizon window — ${MAC}_{H}$
$\operatorname{MAC}_{V}=\frac{1}{H-1}\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|$ (23)
$\operatorname{MAC}_{H}=\frac{1}{H-1}\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|$ (24)
@@ 4.3.1 Stability of Probabilistic Forecasts
To assess stability across time, SDC evaluates both vertical stability (across forecast updates) and horizontal stability (within a forecast horizon). For $N$ number of scenarios and a forecasting horizon $H$, these are defined as:
Vertical Stability (SDCV):
$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (25)
@@
$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (26)
4.3.2 Accuracy of Point Forecasts